# U-net based prediction of cerebrospinal fluid distribution and ventricular reflux grading

### v1

@@ 3.5 Ventricular reflux grading
Traditionally, CSF diagnostics with MRI have primarily focused on anatomical abnormalities in the brain (Ahmad et al., 2021). To analyze cerebrospinal fluid dynamics, a grading scheme was introduced and since then used in clinical assessment of patients (Eide et al., 2020, 2021b): Ventricular reflux refers to the retrograde movement of CSF from the SAS back into the cerebral ventricles and is measured on an integer scale from 0 to 4. Visually assessed reflux can be categorized into five grades based on the intensity and duration of tracer enrichment within the ventricles, which serves as a critical parameter in evaluating fluid dynamics and can be linked to the efficiency of CSF clearance (Eide et al., 2021b):
@@
Reflux grades 3 and 4 have been shown associated with raised intracranial pressure pulsatility, a feature of shunt-responsive iNPH (Eide et al., 2021b, 2020). Comparing MRI biomarkers of CSF tracer dynamics (including ventricular reflux grades) with anatomical MRI biomarkers of CSF space, markers of neurodegeneration and pulsatile intracranial pressure score showed reduced tracer clearance and higher grades of ventricular reflux (grades 3–4 in shunt-responsive idiopathic normal pressure hydrocephalus) (Eide et al., 2020). In this work, we use the ventricular reflux grading scheme as an additional method of model performance evaluation.
@@ 4.1 Image prediction from baseline and tracer-enriched scans
Mean squared error (MSE) and mean absolute error (MAE) of predicted and real scans , each of dimension 256×256×2 (axial and sagittal plane), were computed using PyTorch. We trained for 250 epochs and achieved MSE and MAE for the testing set and MSE and MAE for the training set, indicating overfitting of the model after 50 to 100 epochs. Visual inspection of predictions showed preservation of anatomical characteristics despite some blurriness. Given that the model utilized 129 training scan pairs compared to several thousands in other frameworks (MNIST database: 60,000 training images and 10,000 test images), this reflects outstanding performance and highlights the promising potential of deep learning-based models, particularly in the medical domain, where routinely performed data collection, curation and storage is still rare. A sample test case is depicted in Figure 5.
@@
Table 2 summarizes achieved errors on the test data set for different MR imaging time intervals used as training data. Mean squared errors resulted in the order of magnitude , while mean absolute errors were in the order of magnitude . Remarkably, training solely on data obtained 1-2 hours post-injection showed slightly better performance (MSE ) compared to tracer-enriched MR scans obtained 3-5 hours, 5-7 hours, or 7-9 hours after intrathecal gadobutrol injection (MSEs , , , respectively). We experimentally found that using the -loss generally predicts anatomical structures with more precision compared to an loss function. Two randomly drawn test cases showing real and predicted MR imaging are displayed in Figure 6. Train and test losses during the U-net training process (-loss) are displayed in Appendix 7.
@@
4.2 Ventricular reflux grading performance compared with neuroradiologist assessments
The U-net model trained on scans displaying the tracer distribution within the first two hours after tracer injection was further evaluated with the ventricular reflux grading scheme, which has been used in clinical CSF patient assessment. For this purpose, three human raters (board-certified neuroradiologists) were asked to assign the ventricular reflux grading on 62 sagittal-axial scan pairs that display tracer distribution after 24 hours; half of which were original scans taken at this time (ground truth, GT), while the other half were corresponding predicted images generated by our model. Raters were blinded both for clinical information and whether an image was real or artificial. To avoid bias by memorizing patient-specific characteristics, rating was done in two installments such that at no time human raters had the chance to compare real and artificial scans of the same individual. Furthermore, this procedure was done for both loss functions to assess different model architecture performance. The mean model prediction accuracy over all three raters was for - and -loss. Because of a natural inter-rater variability in ventricular reflux grading (Eide et al., 2020), we further used Krippendorff’s alpha with interval metric to evaluate the model performance. This statistical measure is designed to assess the reliability of coding in content analysis, particularly when multiple observers are involved, and offers high flexibility in handling different types of data, including nominal, ordinal, interval, and ratio scales. Krippendorff’s alpha is defined as the proportion of observed agreement to the agreement expected by chance, adjusted for the level of disagreement among raters. Its mathematical definition is given by , where denotes the observed disagreement, and represents the expected disagreement under the assumption of random coding. This formula ensures that the measure accounts not only for simple agreement but also adjusts for chance agreement, which is crucial in providing an accurate estimate of reliability. Its range extends from negative values, indicating systematic disagreements, to 1, reflecting perfect agreement among coders. An alpha value of 0 denotes agreement equivalent to that expected by chance. In practice, values above 0.8 are typically considered indicative of high reliability, although the threshold for acceptability can vary depending on the context of the study and the nature of the data being analyzed. In this analysis, agreement of three human raters each and the model (mean 0.920, range 0.884 - 0.936 for -loss) evaluated equal or superior compared with inter-rater reliability between three human raters on GT (0.856 for -loss). A full analysis of Krippendorff’s alpha values is shown in table 3.
| U-Net (-loss) | U-Net (-loss) | |
| Rater 1 on GT and predictions | 0.940 | 0.871 |
| Rater 2 on GT and predictions | 0.936 | 0.822 |
| Rater 3 on GT and predictions | 0.884 | 0.869 |
| All raters on GT | 0.856 | 0.869 |
| All raters on predictions | 0.878 | 0.873 |
@@ 5 Discussion
The main observation of this study was that U-net based deep learning provided a high pixel-wise accuracy in predicting the 24-hour distribution imaging utilizing images obtained within 1-2 hours post-injection. These results suggest that the current clinical practice of extended MR imaging sessions to capture later stages of tracer distribution could be reduced without sacrificing diagnostic accuracy. This may provide for improved cost-efficiency and patient throughput.
Traditionally, phase-contrast magnetic resonance imaging (PC-MRI) has been used to assess CSF flow in vivo (Eide et al., 2021b); however, this method has limitations, and tracer studies using contrast agents as CSF tracers may have advantages. A disadvantage with tracer studies is the need for several MRI acquisitions, which not only place significant burden on patients, but are also resource-intensive. The present results document that deep learning-based approaches may be used to avoid repetitive MRI acquisitions.
@@
The machine learning approach may even be better than the human assessment. In addition to error metrics common in neural network regression, we evaluated the model performance by comparing the grading of ventricular reflux on both real scans and those predicted by our model. Inter-rater reliability between the human raters and the model was generally higher than the reliability between human raters alone, indicating that the artificially generated MR scans of the human brain could be equally conclusive for diagnostic CSF assessment.
@@
Some limitation should be noted. While the proposed U-net-based deep learning model demonstrates potential in predicting CSF tracer distribution and assessing ventricular reflux, two limitations of this study should be noted. First, our study is based on a small cohort size of 136 patients with different underlying diseases, which may limit the generalizability of the model’s performance across diverse populations. Second, we utilize a 2D U-net architecture for image prediction, which limits its effectiveness in capturing the full complexity of 3D anatomical structures. Utilization of 3D volumetric data is desirable in the clinical setting, as neuroradiologists commonly assess CSF dynamics and ventricular reflux grading scores across multiple planes.
@@
Future prospects are the use of sufficient computational resources that will give further research opportunities, ideally leveraging volumetric 3D data and integrating this data into a 3D U-net architecture. Another promising direction for future research lies in predicting specifically the tracer clearance process, particularly after the peak distribution at roughly 24 hours post-injection. It remains an open question whether data from the first one to ten hours post-injection are satisfactory to predict decline of tracer signal with a similar deep-learning approach, which would provide further insights into potentially disturbed flow patterns. Understanding the clearance process is crucial, as delayed clearance serves as pathophysiological indicator of several conditions related to neurodegenerative decline, such as in idiopathic normal pressure hydrocephalus (iNPH), Alzheimer’s, and Parkinson’s disease (Hornkjøl et al., 2022; Lopes et al., 2022). Furthermore, integrating clinical measurements, outcomes and longitudinal data, such as patient demographic information, neurocognitive assessment scores, blood work, CSF biomarkers, and patient-specific tracer concentration could enhance the prediction performance and offer more comprehensive evaluations of the model’s predictive capabilities and clinical utility.
@@
In summary, U-net based deep learning was found useful for predicting tracer distribution 24 hours after its intrathecal administration. This may have clinical implications by reducing the number of MRI acquisitions, but further research is required to implement such a methodology in 3D MRI that is used clinically.
@@
Appendix A Appendix
A.1 Loss curves for U-net training and testing over 250 epochs ( loss)

### v2

@@ 3.5 Ventricular reflux grading
Traditionally, CSF diagnostics with MRI have primarily focused on anatomical abnormalities in the brain[47]. To analyze cerebrospinal fluid dynamics, a grading scheme was introduced and since then used in clinical assessment of patients [27]: Ventricular reflux refers to the retrograde movement of CSF from the SAS back into the cerebral ventricles and is measured on an integer scale from 0 to 4. Visually assessed reflux can be categorized into five grades based on the intensity and duration of tracer enrichment within the ventricles, which serves as a critical parameter in evaluating fluid dynamics and can be linked to the efficiency of CSF clearance[23]:
@@
For detailed visual examples of these grades, readers are referred to the original grading studies conducted by Eide et al.[23] Reflux grades 3 and 4 have been shown associated with raised intracranial pressure pulsatility, a feature of shunt-responsive iNPH[23, 27]. Comparing MRI biomarkers of CSF tracer dynamics (including ventricular reflux grades) with anatomical MRI biomarkers of CSF space, markers of neurodegeneration and pulsatile intracranial pressure score showed reduced tracer clearance and higher grades of ventricular reflux (grades 3–4 in shunt-responsive idiopathic normal pressure hydrocephalus) [27]. In this work, we use the ventricular reflux grading scheme as an additional method of model performance evaluation.
@@ 4.1 Image prediction from baseline and tracer-enriched scans
Mean squared error (MSE) and mean absolute error (MAE) of predicted and real scans , each of dimension 256×256×2 (axial and sagittal plane), were computed using PyTorch. We trained for 250 epochs and achieved MSE and MAE for the testing set and MSE and MAE for the training set, indicating overfitting of the model after 50 to 100 epochs. Visual inspection of predictions showed preservation of anatomical characteristics despite some blurriness. Figure 4 presents a sample test case. A comparison between columns 4(c) and 4(e) with 4(b) reveals that the model slightly underestimates tracer enrichment intensity, particularly in the parietal lobe. Furthermore, comparing columns 4(d) and 4(f) with 4(b) demonstrates that the -loss function yields visibly better performance than the -loss function. This pattern was consistently observed across nearly all test cases.
Given that the model utilized 105 2D training MRIs compared to several thousands in other frameworks (MNIST database: 60,000 training images and 10,000 test images), this reflects satisfactory performance and highlights the potential of deep learning-based models to deliver meaningful results even in scenarios where medical imaging datasets are small and less standardized.
@@
Table 2 summarizes achieved errors on the test data set for different MR imaging time intervals used as training data. Mean squared errors resulted in the order of magnitude , while mean absolute errors were in the order of magnitude . Remarkably, training solely on data obtained 1-2 hours post-injection showed performance (MSE ) comparable to the model trained with tracer-enriched MR scans obtained 3-5 hours, 5-7 hours, or 7-9 hours after intrathecal gadobutrol injection (MSEs , , , respectively). Train and test losses during the U-net training process (-loss) are displayed in Appendix 7.
Two randomly drawn test cases showing real and predicted MR imaging are displayed in Figure 5. The T1-weighted MRI scans illustrate two sample test cases (sagittal and axial planes) of CSF tracer (gadobutrol) distribution in the subarachnoid and intraventricular spaces, taken 1–2 hours (first column) and 24 hours (second column) after intrathecal tracer injection. At the first time-point, enhancement with tracer (gadobutrol) induces an increased signal (white), whereas unenhanced CSF is dark. It is interesting to note that in in Patient 1, the sagittal images (upper row) at 1-2 hours after tracer injection show enhancement within an enlarged cisterna magna dorsal to the cerebellum and in the cisternal spaces ventral to the brain stem. The enhancement is well predicted by the neural network. Furthermore, there is yet no ventricular tracer reflux. At 24 hours (second column), the tracer has reached all parts of the ventricular system, though hypointense to the basal cisternal spaces (reflux grade 3). The level of enhancement is well predicted by the network. Ventricular reflux grade 3 is also seen in in Patient 2 (lower two rows), where we also note that the frontal enhancement is well-predicted. In both patients, tracer enhances unevenly in cerebral sulci at the surface. These features are all predicted well as shown in columns 5(c) and 5(e). The involved medical experts found that using the -loss generally predicts anatomical structures with more precision compared to an -loss function. This is confirmed comparing columns 5(d) and 5(f), which show the absolute pixel-wise error of the predictions compared to the target column 5(b). Differences between the two loss functions are most visible in the frontal and parietal lobe.
@@
4.2 Ventricular reflux grading performance compared with neuroradiologists’ assessments
The U-net model trained on scans displaying the tracer distribution within the first two hours after tracer injection was further evaluated with the ventricular reflux grading scheme. For this purpose, three human raters (board-certified neuroradiologists) were asked to assign the ventricular reflux grading on the test set of size Ntest = 31 comprising sagittal-axial scan pairs that display tracer distribution after 24 hours. Grades were assigned to the original scans taken at this time (referred to as ground truth, GT) and corresponding predicted images generated by our model. Raters were blinded both for clinical information and whether an image was real or artificial. Raters were also blinded to pre-injection imaging and evaluated only the scans obtained at the 24h mark. To avoid bias by memorizing patient-specific characteristics, rating was done in two installments such that at no time human raters had the chance to compare real and artificial scans of the same individual. Furthermore, this procedure was done for both loss functions to assess different model architecture performance.
Because of a natural inter-rater variability in ventricular reflux grading[27], we further used grade transition matrices (horizontal transition) to evaluate performance on the test set for each rater, shown in Figure 6. These matrices illustrate the alignment between grades assigned to real axial-sagittal image pairs per subject and their corresponding pairs of model predictions. For instance, Figure 6(a) shows that 92% of individuals in the test set who were assigned a grade of 3 by Rater 1 on the real images also received the same grade 3 from Rater 1 on the predicted images.
Rater 1 exhibited a strong agreement, with most cases aligning along the diagonal, indicating consistency in grading. However, slight discrepancies were observed in lower-grade assignments, where some predicted grades deviated by one category. Rater 2 showed a similar pattern, though with a marginally higher proportion of off-diagonal transitions, particularly for grade 3 cases being over-predicted as grade 4 or grade 4 cases being under-predicted as grade 3. Noticeably, none of the scans that were assigned grade 0 on the real MRI were assigend grade 4 on the artificial MRI for any of the raters, likewise the other way around.
Lastly, Rater 3 demonstrated the highest variability, with 14% of grade 0 cases being assigned grade 3 on the predictions, which provides a contrast to Raters 1 and 2. This behavior could be explained by the fact that the distinction between grade 0 and 3 hinges solely on the presence or absence of signal enhancement at 24h, making the classification potentially ambiguous in regions of interest with low enhancement.
In summary, the findings of this work demonstrate that the model’s predictions lead to consistent grades similar to those that are obtained with the ground truth. Variations observed in borderline cases are likely to reflect differences in individual grading tendencies rather than significant limitations of the model itself. It is worth to point out thtat grades 1 and 2 do not occur because the data used for grading did not exhibit transient enrichment of the lateral ventricles, which is the defining characteristic of these grades. Tracer distribution patterns observed after 24 hours generally corresponded to Grades 3 or 4, where the enhancement is either lasting but less intense than the surrounding subarachnoid spaces (Grade 3) or fully isointense (Grade 4).
@@ 5 Discussion
The main observation of this study was that U-net based deep learning provided a high pixel-wise accuracy in predicting the 24-hour distribution utilizing images obtained within 1-2 hours post-injection. These results suggest that the current clinical practice of extended MR imaging sessions to capture later stages of tracer distribution could be reduced without sacrificing diagnostic accuracy. This may provide for improved cost-efficiency and patient throughput.
Traditionally, phase-contrast magnetic resonance imaging (PC-MRI) has been used to assess CSF flow in vivo[23]; however, that method has limitations by its ability to only image water motions at short term and over short distances, and tracer studies using contrast agents as CSF tracers may have advantages. In particular, traditional PC-MRI only measures back-and-forth movement on the second scale and includes neither netflow nor the slow pulsations of vasomotion, which may be important for solute transfer. Furthermore, the methodology is not well-suited for the thin and complex spaces of the subarachnoid space [31]. Intrathecal contrast enhanced MRI, such as those described in [22, 34], provide a unique view into the slow transport (over days) of solutes within the CSF and the brain. In fact, as shown by [23], neither ICP score nor aquaductal flow, as measured with PC-MR directly, correlates with contrast enhancement. As such, it is remarkable that the U-net approach, learning from just 2D sample images, manages to predict then enhancement pattern the next day to a reasonable accuracy presented above.
A disadvantage with tracer studies is the need for several MRI acquisitions, which not only place significant burden on patients, but are also resource-intensive.
@@
In addition to error metrics common in neural network regression, we evaluated the model performance by comparing the grading of ventricular reflux on both real scans and those predicted by our model. Inter-rater reliability between the human raters and the model was generally higher than the reliability between human raters alone, indicating that the artificially generated MR scans of the human brain could be equally conclusive for diagnostic CSF assessment.
@@
While the proposed U-net-based deep learning model demonstrates potential in predicting CSF tracer distribution and assessing ventricular reflux, two limitations of this study should be noted. First, our study is based on a small cohort size of 136 patients with different underlying diseases, which may limit the generalizability of the model’s performance across diverse populations. Second, we utilize a 2D U-net architecture for image prediction, which limits its effectiveness in capturing the full complexity of 3D anatomical structures. Utilization of 3D volumetric data is desirable in the clinical setting, as neuroradiologists commonly assess CSF dynamics and ventricular reflux grading scores across multiple planes.
@@
Future prospects are the use of sufficient computational resources that will give further research opportunities, ideally leveraging volumetric 3D data and integrating this data into a 3D U-net architecture. Another promising direction for future research lies in predicting specifically the tracer clearance process, particularly after the peak distribution at roughly 24 hours post-injection. It remains an open question whether data from the first one to ten hours post-injection are satisfactory to predict decline of tracer signal with a similar deep-learning approach, which would provide further insights into potentially disturbed flow patterns. Understanding the clearance process is crucial, as delays in this process serve as a pathophysiological indicator of several conditions related to neurodegenerative decline, such as in idiopathic normal pressure hydrocephalus (iNPH), Alzheimer’s, and Parkinson’s disease [48]. Furthermore, integrating clinical measurements, outcomes and longitudinal data, such as patient demographic information, neurocognitive assessment scores, blood work, CSF biomarkers, and patient-specific tracer concentration could enhance the prediction performance and offer more comprehensive evaluations of the model’s predictive capabilities and clinical utility.
@@
In summary, U-net based deep learning was found useful for predicting tracer distribution 24 hours after its intrathecal administration, providing the network with 2D MRI data. This could have clinical implications by reducing the number of MRI acquisitions; however, further research is needed to implement this methodology for clinical use.
Data availability statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@
Appendix A Loss curves for U-net training and testing over 250 epochs ( loss)